OpenStudy (anonymous):

HELP PLEASE!!! ILL GIVE MEDAL OR WHATEVER YOU WANT JUST HELP!!! Its six problems but in one big problem.

OpenStudy (anonymous):

where is it?

OpenStudy (anonymous):

Ill post in just a sec

OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations. graph of line going through ordered pairs (0, 300) and (450, 0)

OpenStudy (anonymous):

u need an advance calculator for this one @CoDsTeR9000

OpenStudy (anonymous):

ill try tho

OpenStudy (anonymous):

cant i need an advance calculator

OpenStudy (anonymous):

@phi help us

OpenStudy (anonymous):

@ganeshie8 help us

OpenStudy (phi):

2x + 3y = 1,200 Change the equation into slope-intercept form. that means "solve for y" I would start by adding -2x to both sides. can you do that ?

OpenStudy (anonymous):

@qwertyboy56 http://web2.0calc.com/

OpenStudy (anonymous):

3y=-2x-1,200

OpenStudy (phi):

now divide both sides by 3 (that means both sides and *all terms*

OpenStudy (anonymous):

y=1 1/2x+400?

OpenStudy (phi):

like this: 2x + 3y = 1,200 add -2x to both sides 2x -2x + 3y = 1200 + -2x (notice the 1200 stays plus) on the left side 2x -2x is 0 and 0 + 3y is 3y 3y = 1200 + -2x or 3y = -2x +1200 now divide by 3 y= (-2/3) x + 1200/3 or y= (-2/3) x + 400

OpenStudy (phi):

match that with y = m x + b m is the slope b is the y-intercept

OpenStudy (anonymous):

Ok so im lost when u get a fraction of (-2/3)

OpenStudy (phi):

ok. here is some background do you remember how to add fractions? example: \[ \frac{1}{5}+ \frac{2}{5} = \frac{1+2}{5} \] does that make sense ?

OpenStudy (anonymous):

Derp... of course it does

OpenStudy (phi):

when you have a common denominator, you can add the numerators (the top numbers) and put the answer over the common denominator.

OpenStudy (phi):

If this is not clear we are in trouble... anyway, the idea works "the other way" \[ \frac{1+2}{5}= \frac{1}{5} +\frac{2}{5} \]

OpenStudy (phi):

for this problem \[ \frac{3y}{3} = \frac{-2x +1200}{3} \] \[ y= \frac{-2x}{3} + \frac{1200}{3} \]

OpenStudy (anonymous):

Ok so quick question do you have kik and if not can u get it? It will make this alot easier. Its an app for txting without needing to buy a plan and its a free app and I see now... wow I cant believe I havent seen that already.... derp

OpenStudy (phi):

no

OpenStudy (anonymous):

Well can you get it cuz u can get it on android and iphone

OpenStudy (phi):

Identify the slope and y-intercept of the equation. that means match the equation with y = m x + b

OpenStudy (anonymous):

Ok im sad but im somewhat lost.....

OpenStudy (phi):

for Describe how you would graph this line using the slope-intercept method. they probably want you to plot the y-intercept (i.e. plot (0,400) ) then use the slope: -2/3 which means "go down 2, and over 3" and plot a new point draw a line through the two points

OpenStudy (phi):

Ok im sad but im somewhat lost..... which question ?

OpenStudy (anonymous):

When I have to find the slope intercept and have to find the y intercept

OpenStudy (phi):

slope-intercept is the name for this type of equation (because it shows the slope and the y-intercept) you have to "pick out" the slope and the y-intercept. to do that you match your equation with the "general form" y = (-3/2) x + 400 y = m x + b m is the "name" people use for the slope b is the name for the y-intercept.

OpenStudy (phi):

when we write y= mx +b we are saying: the number in front of the x is the slope and the number we add is the y-intercept

OpenStudy (anonymous):

Oooohhhh ok ok...... I see it now... but just to be sure imma write down this stuff for future reference..

OpenStudy (anonymous):

Ok but why is there a y where we need to find x and the actual y intercept is a b??

OpenStudy (anonymous):

This always bugged me

OpenStudy (phi):

can you ask the question a different way ?

OpenStudy (anonymous):

Ok so like the y at the start of the equation when you say y= the y is actually meaning tthe x intercept... why? And the b is actually the y intercept.... WTF???

OpenStudy (phi):

the y does not mean the x-intercept Here is the idea: a straight line has (many!) points on it. If we plot any two points, we can draw the line (through those two points) But people wanted a short formula to figure out what points are on a line as we go "sideways" along the x-axis the line will be at some y value. the short-hand formula tells us what y value (for that line) in this problem y= -2/3 x +400 is the formula. when x is 0, we "follow the instructions", multiply 0 times -2/3 (to get 0) and add to 400 to get 0+400= 400 that means the point (0,400) is on the line we can do that for any x value to find the correct y value and figure out what point is on the line

OpenStudy (anonymous):

Ok so two things... first are u a tutor or liike a teacher cuz like u really will take the time to explain the problem out to me and thats not easy and im super grateful and secondly I see now and ur ****ing amazing srry for cussing but common words cannot explain how awesome u r for helping me(im not sucking up fyi)

OpenStudy (phi):

can you figure out the slope and y-intercept for this line ?

OpenStudy (anonymous):

Idk let me try... ;)

OpenStudy (anonymous):

Ok so now can you give a quick example of an easier problem off the top of ur head that shows how to get the answer from here and I think I can get it cuz I learn from one on one teaching and examples so yea can u? If not thats fine

OpenStudy (phi):

this is memorizing when you see y = (-3/2) x + 400 you remember: the slope is the number in front of the x the y-intercept is the number we add

OpenStudy (phi):

y= 2x + 1 slope is 2 y-intercept is 1 y = (1/2) x - 3 slope is 1/2 y-intercept is -3

OpenStudy (anonymous):

Ok so you divided both intergers by 2?